The closing prices for WTI Crude Oil (expressed in U.S. dollars per barrel) over a ten-day period are shown below: 67.54 69.25 70.37 68.59 68.99 68.86 69.75 70.91 70.37 70.78 a) Compute the sample mean, median, and mode for this data. Don’t forget appropriate units. b) Compute the range, sample standard deviation, and sample variance for this data. Don’t forget appropriate units. If necessary, round to 3 decimal places.
Answer #1

putting above data in ascending order and calculating:

(x-x) S.No 67.54 68.59 68.86 68.99 69.25 69.75 70.37 70.37 70.78 70.91 695.41 (X-X) 4.0040 0.9044 0.4638 0.3036 ().缣纏/ 0.0437 0.6872 0.6872 1.5351 1.8742 10.5879 1.176 1.085 1 -2.001 -0.951 -0.681 -0.551 -0.291 0.209 2 4 0.829 1.239 1.369 10 2x 69.541 s-(x-X)/(n-1)-

a)

sample mean=69.541 USD/Barrel

median=(n+1)/2 th value =5.5 th value=(69.25+69.75)/2 =69.50 USD/Barrel

mode =70.37 (as it appears maximum times (twice)) USD/Barrel

b) range =maximum -minimum =70.91-67.54 =3.37 USD/Barrel

sample std deviation =1.176 USD/Barrel

variance =1.085 USD2/Barrel2